OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5 over 9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 10 degrees Celsius is.

OpenStudy (alexwee123):

: C = 5 over 9 (F − 32) 9C/5=F-32 F=9C/5 +32

OpenStudy (alexwee123):

F=9x10/5 +32 F=50

OpenStudy (anonymous):

thank you!!! F=9x10/5 +32 F=50 is the second part right??

OpenStudy (alexwee123):

yep